Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: In A configuration (08 Jun); 4-band dipoles off (all L band projects enabled). Quasi-stable system for testing (full correlator hardware available). Work to enhance allowed data rates. Some local staff returning this week! Double maintenance day Wed-Thu.
  • Key dates:
    • 30 Jun - Call for Proposals
    • 08 Jun - 12 Sep: A configuration (move from 31 May - 08 Jun).
    • 01 Aug - Proposal Deadline (C, CnB, B)
  • Staff Travel:
    • LS: 9 May - 20 Jun
    • GvM: 25 May - 24 Jun
    • CC: 19 Jun - 26 Jun
    • MC: 18 Jun - 26 Jun
    • CC: 28 Jun - 30 Jun (CARMA TAC)
    • RP: 10 Jul - 23 Jul (UK, SA)
    • RP: 09 Aug - 21 Aug (URSI)
    • FO: 12 Aug - 12 Sep
    • RP: 10 Sep - 16 Sep (Brazil)
  • Visitors:
    • Robert Harris (AC982): 24 Apr - 16 Jul (High frequency w/ Claire)
    • Bob Sault: 16 May - 23 Jun (V Stokes; wobbles, splatter and other messy areas)
    • Michiel Brentjens: 11 Jun - 11 Nov (Holography; characterizing polarized beams)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
    • Antennas: EA21 back to array 11 Jul. Report from ACU consultant due.
    • Receivers: EA12 X band; EA21 spate of receivers by the end of the month (28 Jun)
    • Calibration (Perley): Gain compression/expansion still being looked at.
  • WIDAR (Rupen):
    • Notes (Rupen/Sowinski): 3-bit samplers returned to the lab for further work
including the ability to remotely change the bits used - heading back out. BlB hardware looks good. New patch to CRM; assessing test suite for StB. Proposed code for delay rocking is under review.

Software status and updates (Butler)

Commissioning Group Reports and Test Planning

  • NOTE: Please cc me on submitted tests so that it can be assigned the appropriate priority!

Commissioning priorities for the week (McMullin, Chandler)

Configuration move week (work around these activities:
  • Mon
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Tue
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Wed: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Thu: Maintenance day
    • 0630-1700: Maintenance operations
    • 1800-0900: Science operations and testing
  • Fri (not-yet-a-holiday):
    • 0900-1700: Testing/software development time
    • 1800-0900: Science operations and testing
  • Sat-Sun
    • Science operations + planned testing

Issues Affecting Science Observing:

  • Correlator: [Rupen/Carlson w/ Momjian/Sault] V stokes issues.Further data taking; new theories; Bob leaves on Thursday; discussion on Tuesday.
  • Correlator: [Pokorny/Rupen/McMullin] time mislabeling in SDM data (non-synchronous operation between correlator dumps and CBE can require additional 1/2 integration time padding for online flags). This week!
  • Correlator: [Clark/Carlson] Delay rocking across bands. Implementation pending factor-of-2 reconciliation.
  • Array: [Gustaaf] Amplitude wobbles seen on subset of antennas; phase jumps at scan boundaries also seen on subset of antennas (Ka, Q bands only thus far).
  • Array: [All] Zoo of other effects (some data are very good, some still seeing antenna-based IF/Polarization issue).

Testing This week/soon

  • V stokes testing (Sault/Momjian) Discussion this afternoon.
  • CM/CBE: New occurrence of missing scan; Friday crash.
  • Astrometry (Bruthaler): UVW precision tests.
  • Correlator: Rupen: S-band shadow testing
  • Spectral line: Momjian: HI High resolution bandpass testing
  • Array: *; Individual issue sleuthing (focus on phase jumps).

Commissioning Groups

  • Big deadline is end of this month for incorporation into links for Call for Proposals due 1 July.

Old OSRO: Retained as is.
New OSRO:
a. All SBPs get a single baseline board pair (BlBP).
b. All SBPs in a BBP form a continuous, non-overlapping, region in frequency space.
c. The polarization choice is made at the BBP level and applies to all SBPs in that BBP. 
d. All BBPs are 8-bit signals.
e. Integration time: minimum is 1.0 s.  RCT is not the place to consider array configurations in validation logic, though NRAO Defaults can have default setups for D,C,B array cfg where Tint is set to 3.0s.  
New RSRO:
a. Relaxation of the data rate.
b. Variable BW per sub-band
b. Availability of 3-bit samplers for subset of antennas.
    • Adv: Low band (Owen):
      • Notes:
Two cases:
1) The frequency will be interpreted as it currently is in OSRO; that is, it is not a
baseband center but rather is a Sky or Rest frequency that, because all the
SBs form a contiguous range of frequencies, sets the frequency range of
interest.

The software will then do its best to ensure good science; currently, what this
means is that it will attempt to place the baseband center such that
it will minimize the number of 128 MHz filter edges touched. For rest frequencies, it
will take the extra step of offsetting by 0.5*sub-band bandwidth to get the
specified frequency in the center of a sub-band.

The easiest case is not to allow the user control over this (yet!) but
let the software determine whether the 0.5*sub-band bandwidth offset will be negative
or positive.

2) Allow specifying two frequencies; the first is as above while the second specifies the offset from
that 'center' frequency to the line of interest (because it is an offset and the
frequencies contiguous, it also would specify the sub-band used). This offers
more flexibility for those attempting to get other lines within their frequency
range of observation but would not offer any corrections to what the user
inputs (they get what they specify).
    • Adv: Astrometry (Mioduszewski):
      • Notes: Begin testing of UVW precision issues (Brunthaler).
  • Calibrator Models
    • Notes: Status report and Ka/S update: 32 and 37 GHz observations available for 3C138, 3C147, 3C48 (3C286 by the end of the week); currently the information is put on the web (rather than added directly to CASA/AIPS; after A configuration data is added, it will be added directly. S band is not yet available.
  • HPC (Bhatnagar/Owen/Golap):
    • Notes:
  • Imaging (Ott/Golap):
    • Notes:
  • Algorithm (Bhatnagar, Owen):

RSRO/ECSO Reports

RSRO A configuration status: https://safe.nrao.edu/wiki/bin/view/EVLA/2011BRSROStatus

Individual reports (updates)

Misc:

  • RSRO Calendar: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROCalendar
  • RSRO DnC, C config status: https://safe.nrao.edu/wiki/bin/view/EVLA/2010CRSROStatus
  • Test Project Codes == *TRSR*
  • Guidelines for RSRO observing: https://safe.nrao.edu/wiki/bin/view/EVLA/RSROObservingPreparationGuidelines
  • Notes:
    • Current Residents
      • D.J. Saikia (10C-119/AI145)::RSRO-75; 300G(2); x7106-POC: Momjian (Low Frequency)
      • J. Irwin (10C-119AI145)::RSRO-75; 300G(1); x7333 -- POC: Myers (Low Frequency)
      • Peter Hofner (10B-124/AH1011)::RSRO-26; 200G(1); x7496 -- POC: Ott (CASA testing)
      • Sergio Dzib (10B-131); 200G;x7496 -- POC: Mioduszewski (Astrometry)
      • Jacqueline van Gorkom (10C-211/AV324); 200D/x7098 -- POC: Momjian (Low Frequency)
      • Laura Chomiuk (*); 200B/x7497 -- POC: Goss (Bandpass stability)
      • Andreas Brunthaler::200E/x7246 -- POC: Mioduszewski (Astrometry/VLBI)
      • Jane Greaves::(AC982)::RSRO-4; 300G;x7106 -- POC: Chandler (High Frequency Documentation for Workshop)

Demo Science Status

AOB

Misc Actions

  • Old/Low priority
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • Development of WIDAR guidelines document (Michael)
-- JosephMcMullin - 2011-06-18
Topic revision: r2 - 2011-06-21, JosephMcMullin
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback